### v2

@@ 5 Data Availability
The datasets generated and analyzed during the current study are available in this repository: https://depositonce.tu-berlin.de/handle/11303/22195.
6 Code Availability
@@
The underlying code for this study is available in the GitHub repository and can be accessed via this link: https://github.com/mohsen-raoufi/messengers.
Acknowledgments
@@
This study was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2002/1 “Science of Intelligence” – project number 390523135. The funder played no role in the study design, data collection, analysis and interpretation of data, or the writing of this manuscript.
@@
Appendix
Relaxation Time of the DMP
An important property of the DMP to consider for studying the transient behavior of Markov Processes is relaxation time ($\tau_{c}$). The relaxation time is the time that it takes for the system to converge to its stationary state [46]:
$\tau_{c}=\frac{1}{p_{\text{E}}+p_{\text{M}}}.$
Fig. 10 illustrates how the relaxation time is determined by transition probabilities. This figure supplements Fig. 7 and clarifies the change in the optimal lower-left part of the parameter space.
Figure 10: Relaxation time of the DMP
Temporal Development of Transient DMP
As discussed in Sec. 2.4.1, having a non-stationary initial condition of the DMP changes the locally optimal configurations depending on the final time of the simulation. This is shown in Fig. 11 for various snapshots of the opinion precision error at different times.
(a) t=2k
(b) t=4k
(c) t=10k
(d) t=20k
(e) t=40k
Figure 11: Time development of parameter space for precision error of a transient DMP. Here we start the simulation with a collective of all Exploiters. The locally optimal regions shift toward slower DMP configurations over time.
One-dimensional Information landscape
Our results for a similar system in one-dimensional opinion space confirm our results for the 2D system. Among the few modifications needed to fit the model to the 1D scenario, updating the random walk had a significant impact on the performance. We know that for random walking agents, the probability of encountering another agent drops significantly with the dimension size. This can also open new doors to investigating the performance of the Messengers in high-dimensional space. The development of agents’ opinions, when the information landscape is one-dimensional, resembles the classic opinion dynamics with bounded confidence, see Fig. 12.
Figure 12: Temporal development of opinions for 1D information space, a) echo chambers emerge without Messengers, b) Messengers bring consensus back to the collective.
Time-Development of Opinions
The time development of agents’ opinions and positions illustrates how collective forms over time. It also demonstrates that the two dimensions of opinion and position are separate, especially for Messengers whose opinions are independent of their position in space. In Fig. 13, we showed how opinions and positions of agents evolve over time for different pairs of DMP parameters. We used the radial mapping of agents’ positions to convert the 2D positions into a one-dimensional variable.
Figure 13: Temporal development of opinions and positions for different DMP parameters. The first and second rows show the position and opinion of agents, respectively. The third rows illustrate the distribution of collective opinion over time with Messengers as white dots. a) Too many Exploiters (R1), b) a few fast-switching Messengers (R2-b), c) a few slow-switching Messengers (R2), d) many fast-switching Messengers (R2-c), e) a few specialized Messengers with no switching, and e) all Messengers (R3).

### v3

@@ 4 Discussion
The datasets generated and analyzed during the current study are available in this repository [68]:
https://doi.org/10.14279/depositonce-20996.
Code Availability
@@
The underlying code for this study is available in the GitHub repository and can be accessed via this link:
https://github.com/mohsen-raoufi/messengers.
Acknowledgements
@@
This study was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2002/1 “Science of Intelligence” – project number 390523135. The funder played no role in the study design, data collection, analysis and interpretation of data, or the writing of this manuscript. We used OpenAI’s ChatGPT to assist with language refinement, grammar correction, and editing the manuscript. All scientific content, analysis, and interpretation were solely developed by the authors.